TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Created by Main.UnknownUser on 2006-12-05 - 18:25

60A converter tests - Tuesday 5th Dec.

60A tests advancements

Advancement of today:

  • AB-PO performed its tests on C24-26-28-30-32_L8
  • Polarity tests on C22-24_L8
  • TS-MME activities on C12-14_L8
  • AT-MEL activities on cells C12-14L8
  • Positioning of the Cable in sector 7-8 finished at 11:30.

Tomorrow activities:

  • AB-PO activities on cells C34_L8 - C34-32-30-28_R7 no problems for co activities.
  • TS-MME activities on C18-20-22-24-26_L8 no problems for co activities
  • Polarity tests on C28-34_L8 no problems for co activities.

Edit, attach file or add comment

Created by Main.MircoCoccoli on 2006-12-05 - 13:50

Preparation for Cool Down sec78 - 5th Dec 06

PDF RAT_20061205_wslides.pdf:
Minutes with Presentation by D.Bozzini attached

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 3 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 04.12.06 Access & Safety Cable Installation (full sector) (J-C.Guillaume; end by Dec 7th)
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 04.12.06 Re-positioning collars used to fix the WRL to the CLs (all DFBs) (A.Perin-D.Bozzini-F.Millet)
  • 29.11.06 Pumping of insulation vacuum for DFBX
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

-- Main.MircoCoccoli - 05 Dec 2006 - 13:47

Edit, attach file or add comment

Created by Main.MirkoPojer on 2006-12-05 - 12:58

Special RAT on SCT in RR57/UJ56 and ECR application, Dec 5

PDF RAT_20061205.pdf:

PowerPoint RQ4-RQ10-ECR.ppt:

Microsoft Excel Spreadsheet RR57-UJ56-8H-24H-PC_LIST.xls:


Please find enclosed the list of PC which will be powered during the 8h heat run and the 24h heat run. The list of steps to be implemented following the ERC (EDMS No.804188) is also included

Open Issues

  • 22.11 Fire detection – awaiting certification
  • 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-12-01 - JacekSzkutnik
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login